# Coupling of magnetic and lattice collective excitations in the 2D van der Waals antiferromagnet FePS3

### v1

@@ VII Optical Circular Conductivity
Figure 12: Circular optical conductivities $\sigma_{+}$ and $\sigma_{-}$ reconstructed separately for each polarization (Pol 1 and Pol 2), using the combination of full-protocol Faraday rotation and absolute transmission spectra fitted using RefFit software.
Figure 13: Circular optical conductivities $\sigma_{+}$ and $\sigma_{-}$ reconstructed for each polarization (Pol 1 and Pol 2) using Eq. 3, where the Hall conductivity $\sigma_{xy}$ was independently derived for each polarization without averaging.
@@
The main text considers an average of the dielectric constant in order to obtain the optical circular conductivity from the Faraday angle rotation and magneto transmission measurements in the two principal axes. Figures 12 and 13 provide complementary representations of the circular optical conductivity, prior to the averaging process, which help assess the consistency of the analysis presented in the main text (Fig. 10).
@@
Figure 12 shows $\sigma_{+}$ and $\sigma_{-}$ reconstructed separately for the two principal polarizations (noted by the two anglea of the analyzer). In this case, the circular conductivities were obtained by combining the Faraday rotation measured using the full rotation protocol with the absolute transmission spectra, and fitting them directly using the MOKE + VDF procedure in RefFit software.
@@
Figure 13 displays $\sigma_{+}$ and $\sigma_{-}$ for the two polarizations however now derived by using Eq. 3 where $\sigma_{xx}$ and $\sigma_{yy}$ were averaged. Here, the off-diagonal term $\sigma_{xy}$ was determined independently for each analyzer configuration without averaging.
@@
Fig. 10 of the main text shows the result obtained by averaging $\sigma_{xy}$ from the two polarizations, providing a single optical circular conductivity spectrum for FePS3. Comparing Figs. 12 and 13 demonstrates that the overall important features of the circular dichroism that we noted in the main text, including the splitting of the magnon modes and the spectral weight asymmetry, are reproducible across all different analysis procedures. Minor differences between the reconstructions mainly reflect systematic effects such as residual cross-polarization signals and slight misalignment of the sample.

### v2

@@ Appendix B Optical Circular Conductivity in the Birefringence Case
Figure 15: Top panel: Circular optical conductivities $\sigma_{+}$ and $\sigma_{-}$ reconstructed separately for each polarization (Pol 1 and Pol 2), using the combination of full-protocol Faraday rotation and absolute transmission spectra fitted using RefFit software. Bottom panel: Circular optical conductivities $\sigma_{+}$ and $\sigma_{-}$ reconstructed for each polarization (Pol 1 and Pol 2) using Eq. 3, where the Hall conductivity $\sigma_{xy}$ was independently derived for each polarization without averaging.
@@
To further demonstrate the negligible effect of the birefringence on the Faraday rotation, we refer to our analysis of the polarized data. The main text considers an average of the dielectric constant in order to obtain the optical circular conductivity from the Faraday angle rotation and magneto transmission measurements in the two principal axes. Figure 15 provides complementary representations of the circular optical conductivity, prior to the averaging process, which help assess the consistency of the analysis presented in the main text (Fig. 12).
@@
Figure 15 (top) shows $\sigma_{+}$ and $\sigma_{-}$ reconstructed separately for the two principal polarizations (noted by the two angles of the analyzer). In this case, the circular conductivities were obtained by combining the Faraday rotation measured using the full rotation protocol with the absolute transmission spectra, and fitting them directly using the MOKE + VDF procedure in RefFit software.
@@
Figure 15 (bottom) displays $\sigma_{+}$ and $\sigma_{-}$ for the two polarizations however now derived by using Eq. 3 where $\sigma_{xx}$ and $\sigma_{yy}$ were averaged. Here, the off-diagonal term $\sigma_{xy}$ was determined independently for each analyzer configuration without averaging.
@@
Fig. 12 of the main text shows the result obtained by averaging $\sigma_{xy}$ from the two polarizations, providing a single optical circular conductivity spectrum for FePS3. Comparing the different plots in Fig. 15 demonstrates that the overall important features of the circular dichroism that we noted in the main text, including the splitting of the 122  $cm^{-1}$ mode and the spectral weight asymmetry, are reproducible across all different analysis procedures. Minor differences between the reconstructions mainly reflect systematic effects such as residual cross-polarization signals and slight misalignment of the sample.